OpenStudy (anonymous):

Please, anyone, I would appreciate some help. Evelyn purchased a dining room set with a 12-month deferred payment plan. The interest rate on the plan was 25.57%. The dining room set must be paid off within two years after the deferment period. What is Evelyn’s monthly payment after the deferment period if the price of the dining room set was $2,875 and she made no payments during the deferment period? $119.79 $154.28 $170.60 $198.68

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

Okay i think we just need to do it like compound interest: \[\Large A = P(1+r/n)^{nt}\] P is the initial amount, 2875. r is the rate 0.2557. t would be 1 for 1 year. n is 12 assuming it's compounded monthly

OpenStudy (anonymous):

Go on please

OpenStudy (agent0smith):

Ohhh wait we need to find the monthly payment... i dunno how to that, sorry :(

OpenStudy (agent0smith):

Maybe find the amount she'd owe after 2 years (after the deferment) and divide that by 24 months... i'll try it... \[\Large A = 2875(1+0.2557/12)^{12*2} \] which gives A = $4768.63 divide that by 24 for monhtly payment = $198.69

OpenStudy (anonymous):

I just calculated it the exact same way, I think its right :)